# Twists of Superconformal Algebras

### v1

@@ Abstract.
We develop the theory of “conformal twists” of superconformal field theories in dimension 3 to 6, extending the well-known analysis of twists for supersymmetric theories. The conformal twists describe all possible inequivalent choices of a nilpotent element in the superconformal algebra. Such twists can give rise to interesting subalgebras and protected sectors of operators, with the Donaldson–Witten topological field theory and the vertex operator algebras of 4-dimensional $\mathcal{N}{=}\hskip 1.00006pt2$ SCFTs being prominent examples. We work mostly with the complexified superconformal algebras, unless explicitly stated otherwise; real forms of the superconformal algebra may have important physical implications, but we leave these subtleties to future work. To obtain mathematical precision, we explain how to extract vertex algebras and $\mathbb{E}_{n}$ algebras from a twisted superconformal field theory using factorization algebras.
@@ 1. Introduction
In the last two decades, representation theorists have analyzed the space of square-zero odd elements of super Lie algebras (see especially Jenkins [Jenkins]). Of particular relevance to us is the work of Duflo and Serganova [DufloSerganova] (see also [GHSS]), who studied simple super Lie algebras over the complex numbers and described the adjoint orbits of such square-zero odd elements of such algebras. Many superconformal algebras, after complexification, are simple, as we discuss in Section 2. For example, for spacetime dimensions three and six, the superconformal algebras correspond to the super Lie algebras $\mathrm{osp}(n|m)$, which were studied in detail by Gruson [Gruson] prior to Duflo and Serganova. In this paper we explain how these structural mathematical results interface with the role of these super Lie algebras as symmetries of superconformal field theories.
@@ 1.1. What Twisting Does to Superconformal Theories
Notice that all of these questions are totally independent from the specific choice of superconformal theory, and can be applied to draw conclusions in a completely uniform way. As a specific case study, we revisit the results of Beem et al in Section 4.6.
@@ Acknowledgements
We thank Philip Argyres, Mitch Weaver, and Brian Williams for helpful discussions on this topic, and particularly Richard Eager, who offered detailed comments on the manuscript, especially in guiding us through the literature. The National Science Foundation supported O.G. through DMS Grant 2042052. ML is supported in part by the National Research Foundation of Korea (NRF) Grant 2021R1A2C2012350. C.E. and O.G. thank their REU students at Amherst College (Osha Jones and Ziji Zhou) and UMass Amherst (Nikos Orginos, Nathaniel Reid, and Sam Silver) for work on two- and three-dimensional versions of these problems, which gave helpful complementary insights.
@@ 3.3.2. Topological Field Theories
Theorem 3.13 (See Theorem 7.12 and Corollaries 7.13, 7.14).
@@ 4 $k^{2}-4k+14+\delta_{kr}-\delta_{k4}$ $4k$ $k^{2}-8k+14+\delta_{kr}-\delta_{k4}$
4.5. Real Group Orbits: The Case of a Chiral Twist
We now discuss the orbit structure of superconformal twisted theories under the action of real superconformal groups in various signatures. A key ingredient is to recognize that the complex orbits are fiber bundles of a nice kind. We begin by explaining this idea, and its power for studying real orbits, in the simplest example before gesturing at some general machinery. We stick to the chiral twists as the situation is already complicated there.
@@ 4.5. Real Group Orbits: The Case of a Chiral Twist
    In signature $(4,0)$ we have $G_{\mathbb{R}}\cap\mathrm{SL}(4,\mathbb{C})\cong\mathrm{SL}(2,\mathbb{H})$, and $G_{\mathbb{R}}\cap\mathrm{SL}(2,\mathbb{C})\cong\mathrm{SU}(2)$. The action of $\mathrm{SU}(2)$ on $\mathbb{CP}^{1}$ is transitive, as is the action of $\mathrm{SL}(2,\mathbb{H})$ on $\mathbb{C}^{4}\backslash\{0\}\cong\mathbb{H}^{2}\backslash\{0\}$. In this case the real orbit space is a single point.
@@
    In signature $(3,1)$ we have $G_{\mathbb{R}}\cap\mathrm{SL}(4,\mathbb{C})\cong\mathrm{SU}(2,2)$, and $G_{\mathbb{R}}\cap\mathrm{SL}(2,\mathbb{C})\cong\mathrm{SU}(2)$. The action of $\mathrm{SU}(2)$ on $\mathbb{CP}^{1}$ is transitive, but the action of $\mathrm{SL}(2,\mathbb{H})$ on $\mathbb{C}^{4}\backslash\{0\}$ is more interesting: there is a continuous family of orbits parameterized by $\mathbb{R}^{\times}$ (corresponding to non-zero values of the split signature Hermitian form on $\mathbb{C}^{2,2}$), but there is a family of orbits with a finite non-Hausdorff locus (corresponding to the null locus for the Hermitian form). (That is, the quotient space by the $\mathrm{SL}(2,\mathbb{H})$-action is non-Hausdorff.)
@@
In principle one can work out these orbit decompositions out by hand, but there are quite general results of Fels–Huckleberry–Wolf (going back to Wolf’s work from the 1960s) that help here, notably in identifying the lower right corner of the diagram.
@@
Let $G$ be a real form of $\mathrm{SL}(n,\mathbb{C})$ of real rank $\rho$. Let $P$ be a parabolic subgroup of $\mathrm{SL}(n,\mathbb{C})$.
@@
  • For general partial flag varieties, an upper bound holds: there are at most $k|W|$ orbits, where $k$ is the number of conjugacy classes of Cartan subgroups in $G$.
@@
$D_{p,4-p}=\mathrm{Gr}(4-r,4)/G^{\text{base}}_{p,4-p}$
is a finite set with some interesting connected but non-Hausdorff topology. After identifying this finite set, we then need to understand the fibers of the map
$\mathcal{O}(4|k)_{r,0}/G_{p,4-p}\to\mathrm{Gr}(4-r,4)/G^{\text{base}}_{p,4-p}=D_{p,4-p}$
@@
$\mathcal{O}(4|k)_{r,0}/G_{p,4-p}\cong D_{p,4-p}\times(\mathrm{GL}(r_{+},\mathbb{C})/G^{\text{fib}}_{p,4-p})$
@@
The number $k$ appearing in the upper bound for the number of orbits $|D_{*,*}|$ is computed for general real forms of $\mathrm{SL}(n,\mathbb{C})$ by Sugiura [Sugiura]. In particular we find $k=1$ in Euclidean signature and $k=3$ in split signature. The remaining Lorentzian signature case is a hermitian symmetric space so we obtain the exact number of orbits as being 6.
We now turn to discussing each signature in turn.
@@ 4.5.1. Signature $(4,0)$
where $\mathbb{H}$ denotes the quaternions. In general, when we write $\mathrm{SL}(n|m,\mathbb{H})$, it refers to the set of quaternionic matrices such that
@@
where $\mathrm{Nm}(q)\in\mathbb{R}_{\geq 0}$ denotes the norm of a quaternion $q$. In particular $\dim_{\mathbb{R}}\mathrm{SL}(n,\mathbb{H})=4n^{2}-1$. In particular, the even part of this superconformal group can be identified with
@@
The real rank of $\mathrm{SL}(2,\mathbb{H})$ is one, so in a partial flag variety of the form $\mathrm{SL}(4,\mathbb{C})/P$, there are always precisely three orbits for the left-action of $\mathrm{SL}(2,\mathbb{H})$.
@@
Consider the $\mathcal{N}=2\ell$ superconformal algebra with $2\ell\neq 4$. In the space of chiral supercharges of rank $(r_{+},0)$, the space of orbits in signature $(4,0)$ can be identified with
$(\mathrm{GL}(2\ell,\mathbb{C})/(\mathrm{GL}(\ell,\mathbb{H}))\times D$
for a finite topological space $D$ with $|D|\leq 12$.
@@
$(\mathbb{R}^{3}\times S^{1})^{\sqcup m}.$
@@ 4.5.2. Signature $(3,1)$
$\mathbb{C}^{2,0},\mathbb{C}^{1,1},\mathbb{C}^{0,2},\mathbb{C}_{\mathrm{null}}\times\mathbb{C}^{1,0},\mathbb{C}_{\mathrm{null}}\times\mathbb{C}^{0,1},(\mathbb{C}_{\mathrm{null}})^{2},$
where $\mathbb{C}_{\mathrm{null}}$ refers to $\mathbb{C}$ with a trivial, completely degenerate metric.
@@ 4.5.3. Signature $(2,2)$
Thus, overall,the set of orbits in this split signature case is identified with $\mathbb{R}^{2}\times(S^{1})^{2}\times D$ for a finite topological space $D$.
4.6. Case Study: 4d $\mathcal{N}=2$ Chiral Supercharges
In this section we will analyze the specific example considered by Beem et al [Beem_2015] in their derivation of vertex algebras from $\mathcal{N}=2$ 4d superconformal field theory. Beem et al are concerned with a chiral supercharge of rank $(2,0)$ that they dub the twisted Schur operators. The analysis above tells us that for this example
$\dim\mathfrak{z}^{\mathbb{C}}_{\mathcal{Q}}$ $=11,$
$\dim\mathfrak{b}^{\mathbb{C}}_{\mathcal{Q}}$ $=8.$
Let us discuss such chiral supercharges in terms of the usual physics notation. We choose an affine patch $\mathbb{R}^{p,4-p}\hookrightarrow C(\mathbb{R}^{p,4-p})$, associated to which there is an embedding $\mathrm{ISO}(p,4-p)\hookrightarrow\mathrm{Conf}(p,4-p)$. Note that this embedding is non-canonical and so the decomposition that we will discuss below is also non-canonical.
The summands of the odd part of the superconformal algebra decompose as $\mathfrak{so}(4,\mathbb{C})\subseteq\mathfrak{so}(6,\mathbb{C})$-representations as follows. If $V_{4}\otimes W_{2}$ is the positive chirality piece of $\mathfrak{sconf}(4|2,\mathbb{C})_{1}$, we have the decomposition

### v2

@@ Abstract.
We develop the theory of “conformal twists” of superconformal field theories in dimension 3 to 6, extending the well-known analysis of twists for supersymmetric theories. The conformal twists describe all possible inequivalent choices of a nilpotent element in the superconformal algebra. Such twists can give rise to interesting subalgebras and protected sectors of operators, with the Donaldson–Witten topological field theory and the vertex operator algebras of 4-dimensional $\mathcal{N}{=}\hskip 1.00006pt2$ SCFTs being prominent examples. We work mostly with the complexified superconformal algebras, unless explicitly stated otherwise; real forms of the superconformal algebra may have important physical implications, but we only discuss these subtleties in a few special cases. To obtain mathematical precision, we explain how to extract vertex algebras and $\mathbb{E}_{n}$ algebras from a twisted superconformal field theory using factorization algebras.
@@ 1. Introduction
In the last two decades, representation theorists have analyzed the space of square-zero odd elements of super Lie algebras (see especially the work of Jenkins and Nakano [JenkinsNakano]). Of particular relevance to us is the work of Duflo and Serganova [DufloSerganova] (see also [GHSS]), who studied simple super Lie algebras over the complex numbers and described the adjoint orbits of such square-zero odd elements of such algebras. Many superconformal algebras, after complexification, are simple, as we discuss in Section 2. For example, for spacetime dimensions three and six, the superconformal algebras correspond to the super Lie algebras $\mathrm{osp}(n|m)$, which were studied in detail by Gruson [Gruson] prior to Duflo and Serganova. In this paper we explain how these structural mathematical results interface with the role of these super Lie algebras as symmetries of superconformal field theories.
@@ 1.1. What Twisting Does to Superconformal Theories
Notice that all of these questions are totally independent from the specific choice of superconformal theory, and can be applied to draw conclusions in a completely uniform way. As a specific case study, we revisit the results of Beem et al in Section 5.3 and relate the approach of Saberi and Williams to ours in Section 8.4.
@@ Acknowledgements
We thank Philip Argyres, Ingmar Saberi, Mitch Weaver, and Brian Williams for helpful discussions on this topic, and particularly Richard Eager, who offered detailed comments on the manuscript, especially in guiding us through the literature. The National Science Foundation supported OG through DMS Grant 2042052. ML is supported in part by the National Research Foundation of Korea (NRF) Grant 2021R1A2C2012350. CE and OG thank their REU students at Amherst College (Osha Jones and Ziji Zhou) and UMass Amherst (Nikos Orginos, Nathaniel Reid, and Sam Silver) for work on two- and three-dimensional versions of these problems, which gave helpful complementary insights. OG completed revisions on this paper at the Max Planck Institute of Mathematics during a lovely and productive visit there, and he acknowledges its support with gratitude.
@@ 3.3.2. Topological Field Theories
Theorem 3.13 (See Theorem 8.12 and Corollaries 8.13, 8.14).
@@ 4 $k^{2}-4k+14+\delta_{kr}-\delta_{k4}$ $4k$ $k^{2}-8k+14+\delta_{kr}-\delta_{k4}$
5. Chiral Twists in 4d Superconformal Algebras: Some Case by Case Analysis
In this section we discuss the orbit structure of superconformal twists under the action of real superconformal groups in various signatures. We stick to the chiral twists as the situation is already complicated there, and we eventually specialize to the situation that Beem et al [Beem_2015] showed to be strikingly rich.
5.1. Real Orbits of Chiral Twists: General Features
A key ingredient is to recognize that the complex orbits are fiber bundles of a nice kind, where the base is often a familiar homogeneous space and the fibers admit clean descriptions via linear algebra. We begin by explaining this idea, and its power for studying real orbits, in the simplest example before gesturing at some general machinery.
@@ 5.1. Real Orbits of Chiral Twists: General Features
    In signature $(4,0)$ we have $G_{\mathbb{R}}\cap\mathrm{SL}(4,\mathbb{C})\cong\mathrm{SL}(2,\mathbb{H})$, and $G_{\mathbb{R}}\cap\mathrm{SL}(2,\mathbb{C})\cong\mathrm{SU}(2)$. The action of $\mathrm{SU}(2)$ on $\mathbb{CP}^{1}$ is transitive, as is the action of $\mathrm{SL}(2,\mathbb{H})$ on $\mathbb{C}^{4}\backslash\{0\}\cong\mathbb{H}^{2}\backslash\{0\}$. In this case the real orbit space is a single point. (That is, there is a single orbit; all nilpotent supercharges are equivalent.)
@@
    In signature $(3,1)$ we have $G_{\mathbb{R}}\cap\mathrm{SL}(4,\mathbb{C})\cong\mathrm{SU}(2,2)$, and $G_{\mathbb{R}}\cap\mathrm{SL}(2,\mathbb{C})\cong\mathrm{SU}(2)$. The action of $\mathrm{SU}(2)$ on $\mathbb{CP}^{1}$ is transitive, but the action of $\mathrm{SL}(2,\mathbb{H})$ on $\mathbb{C}^{4}\backslash\{0\}$ is more interesting: there is a continuous family of orbits parameterized by $\mathbb{R}^{\times}$ (corresponding to non-zero values of the split signature Hermitian form on $\mathbb{C}^{2,2}$), but there is also a family of orbits with a finite non-Hausdorff locus (corresponding to the null locus for the Hermitian form).
    That is, the quotient space by the $\mathrm{SL}(2,\mathbb{H})$-action is non-Hausdorff.
@@
In principle one can work out these orbit decompositions by hand, but there are quite general results of Fels–Huckleberry–Wolf (going back to Wolf’s work from the 1960s) that help here, notably in identifying the lower right corner of the diagram.
@@
Let $G$ be a real form of $\mathrm{SL}(n,\mathbb{C})$ of real rank $\rho$. Let $P$ be a parabolic subgroup of $\mathrm{SL}(n,\mathbb{C})$. Then the following hold:
@@
  • For general partial flag varieties, an upper bound holds: there are at most $c|W|$ orbits, where $c$ is the number of conjugacy classes of Cartan subgroups in $G$.
@@
$D^{(r)}_{p,4-p}=\mathrm{Gr}(4-r,4)/G^{\text{base}}_{p,4-p}$
is a finite set with some interesting topology that is connected but non-Hausdorff. After identifying this finite set, we then need to understand the fibers of the map
$\mathcal{O}(4|k)_{r,0}/G_{p,4-p}\to\mathrm{Gr}(4-r,4)/G^{\text{base}}_{p,4-p}=D^{(r)}_{p,4-p}$
@@
$\mathcal{O}(4|k)_{r,0}/G_{p,4-p}\cong D^{(r)}_{p,4-p}\times(\mathrm{GL}(r,\mathbb{C})/G^{\text{fib}}_{p,4-p})$
@@
The number $c$ appearing in the upper bound for the number of orbits $|D^{(r)}_{*,*}|$ is computed for general real forms of $\mathrm{SL}(n,\mathbb{C})$ by Sugiura [Sugiura]. In particular we find $c=1$ in Euclidean signature and $c=3$ in split signature. The remaining Lorentzian signature case is a hermitian symmetric space so we obtain the exact number of orbits as being 6.
5.2. The Case $r=2$
We now turn to discussing $r=2$ for each signature in turn.
@@ 5.2.1. Signature $(4,0)$
where $\mathbb{H}$ denotes the quaternions. (Note that for Euclidean signature, we must have $k$ even. Thus take $k=2\ell$.) In general, when we write $\mathrm{SL}(n|m,\mathbb{H})$, it refers to the set of quaternionic matrices such that
@@
where $\mathrm{Nm}(q)\in\mathbb{R}_{\geq 0}$ denotes the norm of a quaternion $q$. In particular $\dim_{\mathbb{R}}\mathrm{SL}(n,\mathbb{H})=4n^{2}-1$.
The even part of this superconformal group can be identified with
@@
We begin by examining the action of $\mathrm{SL}(2,\mathbb{H})$ on the base space $\mathrm{Gr}(2;4;\mathbb{C})$.
Lemma 5.7.
There are two orbits for the action of $\mathrm{SL}(2,\mathbb{H})$ on the base space $\mathrm{Gr}(2;4;\mathbb{C})$. That is,
$D^{(2)}_{4,0}=\mathrm{Gr}(2,4;\mathbb{C})/\mathrm{SL}(2,\mathbb{H})$
has two elements.
Proof.
As the quaternions $\mathbb{H}$ forget to a copy of $\mathbb{C}^{2}=\mathbb{C}\oplus j\mathbb{C}$, we can identify
$\mathbb{C}^{4}\cong\mathbb{C}^{2}\otimes_{\mathbb{C}}\mathbb{H}\cong\mathbb{H}^{2}.$
In explicit terms, equip $\mathbb{C}^{2}$ with the basis $\{e_{1},e_{2}\}$ and let $\mathbb{H}^{2}$ have the same basis (under extension of scalars). Then $\mathbb{C}^{4}$ has a $\mathbb{C}$-linear basis by $\{e_{1},je_{1},e_{2},je_{2}\}$.
Given a subspace $V\subset\mathbb{C}^{4}$ with $\dim_{\mathbb{C}}(V)=2$, representing a point $[V]\in\mathrm{Gr}(2,4)$, we obtain an $\mathbb{H}$-linear subspace $V_{\mathbb{H}}\subset\mathbb{H}^{2}$ by taking the $\mathbb{H}$-span of $V$, i.e.,
$V_{\mathbb{H}}=V\otimes_{\mathbb{C}}\mathbb{H}.$
If $V$ has complex basis $\{v_{1},v_{2}\}$, then
$V_{\mathbb{H}}=\mathbb{H}v_{1}\oplus\mathbb{H}v_{2}\subset\mathbb{H}^{2}.$
This subspace has
$0<\dim_{\mathbb{H}}(V_{\mathbb{H}})\leq 2,$
namely either 1 or 2.
Both values 1 and 2 are realized in practice: the complex subspace
$\mathbb{H}e_{1}=\mathbb{C}e_{1}\oplus\mathbb{C}(je_{1})\subset\mathbb{C}^{4}$
has $\dim_{\mathbb{H}}(\mathbb{H}e_{1})=1$, while the subspace
$W=\mathbb{C}e_{1}\oplus\mathbb{C}e_{2}\subset\mathbb{C}^{4}$
has $\dim_{\mathbb{H}}(W_{\mathbb{H}})=2$. Thus the function $\mathrm{Gr}(2,4)\to\{1,2\}$ sending $V$ to $\dim_{\mathbb{H}}(V_{\mathbb{H}})$ is surjective.
It remains to show that the action of $\mathrm{SL}(2,\mathbb{H})$ is transitive on the fibers of this surjection, as this will show that there are two orbits.
The subspaces with $\dim_{\mathbb{H}}(V_{\mathbb{H}})=1$ parametrize $\mathbb{H}\mathbb{P}^{1}$, so we have an inclusion
$\mathbb{H}\mathbb{P}^{1}\subset\mathrm{Gr}(2,4)$
and $\mathrm{SL}(2,\mathbb{H})$ acts transitively on $\mathbb{H}\mathbb{P}^{1}$ (i.e., every such subspace is of the form $A(\mathbb{H}e_{1})$ for some $A\in\mathrm{SL}(2,\mathbb{H})$).
On the other hand, if $\dim_{\mathbb{H}}(V_{\mathbb{H}})=2$, there is a complex basis $\{v_{1},v_{2}\}$ and these span $\mathbb{H}^{2}$ over the quaternions. Thus, there is some $A\in\mathrm{GL}(2,\mathbb{H})$ such that $Ae_{j}=v_{j}$. This matrix $A$ might not have determinant of norm 1, but by rescaling the basis vectors to get a new basis $\{v^{\prime}_{1},v^{\prime}_{2}\}$, we can find $A^{\prime}\in\mathrm{SL}(2,\mathbb{H})$ such that $A^{\prime}e_{j}=v^{\prime}_{j}$. In other words, $\mathrm{SL}(2,\mathbb{H})$ is transitive on this fiber. ∎
@@
Consider the $\mathcal{N}=2\ell$ superconformal algebra with $2\ell\neq 4$. In the space of chiral supercharges of rank $(r,0)$, the space of orbits in signature $(4,0)$ can be identified with
$(\mathrm{GL}(2\ell,\mathbb{C})/\mathrm{GL}(\ell,\mathbb{H}))\times\{1,2\}$
where $\{1,2\}$ labels the $\mathbb{H}$-dimension of the $\mathbb{H}$-span in $\mathrm{Gr}(2,4)$.
@@
$(\mathbb{R}^{3}\times S^{1})\times\{1,2\}$
where $\{1,2\}$ labels the $\mathbb{H}$-dimension of the $\mathbb{H}$-span in $\mathrm{Gr}(2,4)$.
@@ 5.2.2. Signature $(3,1)$
$\mathbb{C}^{2,0},\mathbb{C}^{1,1},\mathbb{C}^{0,2},\mathbb{C}_{\mathrm{null}}\times\mathbb{C}^{1,0},\mathbb{C}_{\mathrm{null}}\times\mathbb{C}^{0,1},\text{and }(\mathbb{C}_{\mathrm{null}})^{2},$
where $\mathbb{C}_{\mathrm{null}}$ refers to $\mathbb{C}$ with the zero quadratic form. That is, the space $D^{(2)}_{3,1}$ is a finite topological space with six points.
Example 5.11.
If $k=2$, we can again analyze the fiber:
$\mathrm{GL}(2,\mathbb{C})/\mathrm{U}(2)\cong\mathrm{SL}(2,\mathbb{C})/\mathrm{SU}(2)\times\mathbb{C}^{\times}/\mathbb{R}^{\times}\cong\mathfrak{h}^{3}\times S^{1}.$
Thus again, topologically, the space of orbits of rank $(2,0)$ chiral supercharges in the $\mathcal{N}=2$ superconformal algebra, in Lorentzian signature, can be identified with
$(\mathbb{R}^{3}\times S^{1})\times D^{(2)}_{3,1}.$
@@ 5.2.3. Signature $(2,2)$
We begin by examining the action of $\mathrm{SL}(4,\mathbb{R})$ on the base space $\mathrm{Gr}(2;4;\mathbb{C})$.
Lemma 5.12.
There are three orbits for the action of $\mathrm{SL}(4,\mathbb{R})$ on the base space $\mathrm{Gr}(2;4;\mathbb{C})$: for $V\subset\mathbb{C}^{4}$ with $\dim_{\mathbb{C}}(V)=2$, its orbit $\mathrm{SL}(4,\mathbb{R})V$ is characterized by
$\dim_{\mathbb{R}}(V\cap\mathbb{R}^{4})\in\{0,1,2\}.$
That is,
$D^{(2)}_{2,2}=\mathrm{Gr}(2,4)/\mathrm{SL}(4,\mathbb{R})$
has three elements.
Proof.
Let $\mathbb{R}^{4}$ denote the real subspace of $\mathbb{C}^{4}$, and let $i\mathbb{R}^{4}$ denote the real subspace of purely imaginary vectors.
Let $V\subset\mathbb{C}^{4}$ with $\dim_{\mathbb{C}}(V)=2$, and hence represent a point $[V]\in\mathrm{Gr}(2,4)$. We will study its intersection with $\mathbb{R}^{4}$ (the “real intersection”).
Observe that the action of $\mathrm{SL}(4,\mathbb{R})$ on $\mathbb{C}^{4}$ is complex-linear, preserves $\mathbb{R}^{4}$ in $\mathbb{C}^{4}$, and commutes with conjugation. Hence this action preserves $\dim_{\mathbb{R}}(V\cap\mathbb{R}^{4})$.
Now we claim
$\dim_{\mathbb{R}}(V\cap R^{4})\leq 2$
because the intersection consists of the conjugation-fixed points in $V$. Let $V^{\rm conj}\subset V$ consist of $v\in V$ such that $\overline{v}=v$. Note that if $v\in V^{\rm conj}$, then $iv$ is in $V$ but $iv$ is not fixed by conjugation. Hence $V^{\rm conj}$ and $iV^{\rm conj}$ are transverse real subspaces of $V$, and hence
$\dim_{\mathbb{R}}(V^{\rm conj}\oplus iV^{\rm conj})=2\dim_{\mathbb{R}}(V^{\rm conj})\leq\dim_{\mathbb{R}}(V)=4,$
yielding our dimension bound.
Notice that there is a subspace $V$ whose intersection has dimension 2, 1, or 0:
  • (2)
    Take $V=\mathbb{C}e_{1}\oplus\mathbb{C}e_{3}$.
  • (1)
    Take $V=\mathbb{C}e_{1}\oplus\mathbb{C}(e_{3}+ie_{4})$ because a linear combination $ze_{1}+w(e_{3}+ie_{4})$ is real if and only if $z$ real, $w$ real, and $iw$ is real (i.e., $w=0$).
  • (0)
    Take $V=\mathbb{C}(e_{1}+ie_{2})\oplus\mathbb{C}(e_{3}+ie_{4})$ because a linear combination $z(e_{1}+ie_{2})+w(e_{3}+ie_{4})$ is real if and only if $z$ real, $w$ real, $iz$ is real, and $iw$ is real (i.e., $z=0=w$).
Hence we know the orbits surject onto the set $\{0,1,2\}$ under the map $V\mapsto\dim_{\mathbb{R}}(V\cap\mathbb{R}^{4})$.
We now explain why the action of $\mathrm{SL}(4,\mathbb{R})$ is transitve on each fiber of this map.
In case the dimension of intersection is 2, then $V\cap\mathbb{R}^{4}$ admits a real basis with two elements $\{v_{1},v_{2}\}$. There is some element $A$ of $\mathrm{SL}(4,\mathbb{R})$ such that $Ae_{1}=v_{1}$ and $Ae_{3}=v_{2}$. (This does not uniquely determine $A$, of course.) Thus $V$ is in the orbit of $\mathbb{C}e_{1}\oplus\mathbb{C}e_{3}$.
In case the dimension of intersection is 1, then $V\cap\mathbb{R}^{4}$ admits a real basis with one element $\{v_{1}\}$. Extend to a complex basis $\{v_{1},v_{2}\}$ where $v_{2}=x+iy$ with $x,y\in\mathbb{R}^{4}$ and $y\neq 0$. We want $A\in\mathrm{SL}(4,\mathbb{R})$ such that $Ae_{1}=v_{1}$ and
$A(e_{3}+ie_{4})=x+iy,$
i.e., $Ae_{3}=x$ and $Ae_{4}=y$. But we know $v_{1},x,y$ are linearly independent over $\mathbb{R}$ so such an $A$ exists.
In case the dimension of intersection is 0, then $V$ has a complex basis $\{v_{1},v_{2}\}$ where $v_{j}=x_{j}+iy_{j}$ with $y_{1}\neq 0$ and $y_{2}\neq 0$. We want $A\in\mathrm{SL}(4,\mathbb{R})$ such that
$A(e_{1}+ie_{2})=x_{1}+iy_{1}$
and
$A(e_{3}+ie_{4})=x_{2}+iy_{2}.$
These vectors $\{x_{1},x_{2},y_{1},y_{2}\}$ are linearly independent over $\mathbb{R}^{4}$, so we can find a matrix $B$ in $\mathrm{GL}(4,\mathbb{R})$ satisfying our condition: take $x_{1}$ to be the first column of $B$, $y_{1}$ to be the second column, and so on. This matrix is $B$ might have $\det(B)\neq 1$, however. This issue is easily fixed by adjusting the basis $\{v_{1},v_{2}\}$ by rescaling. ∎
@@
Thus, overall,the set of orbits in this split signature case is identified with
$\mathbb{R}^{2}\times(S^{1})^{2}\times\{0,1,2\}$
where the finite set identifies the orbit of the $\mathrm{SL}(4,\mathbb{R})$ action.
5.3. The Case of 4d $\mathcal{N}=2$ Chiral Supercharges
In this section we will analyze the specific example considered by Beem et al [Beem_2015] in their derivation of vertex algebras from 4d $\mathcal{N}=2$ superconformal field theory. Beem et al are concerned with a chiral supercharge of rank $(2,0)$ that they dub the twisted Schur supercharge. The analysis above tells us that for this example
$\dim\mathfrak{z}^{\mathbb{C}}_{\mathcal{Q}}=11\quad\text{and}\quad\dim\mathfrak{b}^{\mathbb{C}}_{\mathcal{Q}}=8.$
Let us discuss such chiral supercharges in the usual physics notation.
We choose an affine patch $\mathbb{R}^{p,4-p}\hookrightarrow C(\mathbb{R}^{p,4-p})$, associated to which there is an embedding $\mathrm{ISO}(p,4-p)\hookrightarrow\mathrm{Conf}(p,4-p)$. Note that this embedding is non-canonical and so the decomposition that we will discuss below is also non-canonical. The summands of the odd part of the superconformal algebra then decompose as representations of $\mathfrak{so}(4,\mathbb{C})\subseteq\mathfrak{so}(6,\mathbb{C})$ as follows. If $V_{4}\otimes W_{2}$ is the positive chirality piece of $\mathfrak{sconf}(4|2,\mathbb{C})_{1}$, we have the decomposition